# Non-commutative Iwasawa theory of abelian varieties over global function fields

### v1

@@ 1. Introduction
It is worth noting that a global function field $F$ with $\mathrm{char}(F)\neq p$ has only one ${\mathbb{Z}}_{p}$-extension, the cyclotomic one (see [28, Proposition 10.3.20]). Therefore, any admissible extension of dimension at least $2$ is necessarily non-commutative. This underscores the relevance of non-commutative Iwasawa theory in the context of function fields. We also note that when $\mathrm{char}(F)=p$, there are infinitely many geometric (i.e., not containing any constant field extension) ${\mathbb{Z}}_{p}$-extensions of $F$ (see [14, Theorem 3]). This presents another intriguing difference between number fields and function fields.
We now return to the core of the paper. Our first main result is as follows:
Theorem 1.3.
Let $A$ be an abelian variety defined over a global function field $F$ with characteristic prime to $p$. Let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Then, the $\mathfrak{M}_{H}(G)$-conjecture holds for the Pontryagin dual $X(A/F_{\infty})$ of the Selmer group ${\mathrm{Sel}}(A/F_{\infty})$.
Furthermore, the Pontryagin dual $X(A/F^{\mathrm{cyc}})$ of the Selmer group ${\mathrm{Sel}}(A/F^{\mathrm{cyc}})$ has $\mu$-invariant equal to zero.
In particular, our result completely resolves Mazur’s conjecture and the $\mathfrak{M}_{H}(G)$-conjecture in the function field setting. This presents a stark contrast with the number field case. It is long established that the Pontryagin dual of the Selmer group over the cyclotomic ${\mathbb{Z}}_{p}$-extension of a number field can have a positive $\mu$-invariant (see [24, §10, Example 2], [13]). In contrast, our result shows that this never occurs in function fields. We note that Sechi [31] has given a proof of this theorem for elliptic curves $E$ without complex multiplication defined over $F$, where $F_{\infty}$ is obtained by adjoining the $p$-power torsion points of $E$ to $F$, under the assumption that ${\mathrm{Sel}}(E/F)$ is finite. In the case $\mathrm{char}(F)=p$, the ${\mathfrak{M}}_{H}(G)$-conjecture has been proven by Ochiai and Trihan under a $\mu=0$ assumption (see [29, Theorem 1.9]). Our proof of this theorem differs entirely from previous work on the ${\mathfrak{M}}_{H}(G)$-conjecture. In particular, Theorem 1.3 provides the first example where the ${\mathfrak{M}}_{H}(G)$-conjecture holds unconditionally. A key element of our proof is the observation that the cohomological dimension of ${\mathrm{Gal}}(\bar{F}/F^{\mathrm{cyc}})$ is equal to one. From there, our approach is essentially cohomological in nature. The authors find it curious that this line of approach seems to have been overlooked in literature.
In light of this result on the $\mathfrak{M}_{H}(G)$-conjecture, we can now attach a characteristic element $\xi_{A}$ to $X(A/F_{\infty})$ in the sense of Coates et al. This allows us to evaluate $\xi_{A}$ at a given Artin representation $\rho:G\to{\mathrm{GL}}_{n}(\mathcal{O})$, where $\mathcal{O}$ is the ring of integers of a finite extension of ${\mathbb{Q}}_{p}$ in ¯$\overline{{\mathbb{Q}}}_{p}$, via the homomorphism $\Phi_{\rho}$ (for details, see Section 7, in particular (7.2)). This yields an element in the field of fractions of the formal power series ring $\mathcal{O}[[T]]$ in an indeterminate $T$ with coefficients in $\mathcal{O}$. For subsequent discussion, we shall introduce the following hypothesis to facilitate explicit calculations:
Hypothesis ($\mathbf{G}$). $H^{i}(G,A_{p^{\infty}}(F_{\infty}))$ is finite for all $i\geq 1$.
We are now in a position to state our next result, presented here in a slightly simplified form. For a more precise and general statement of the theorem below, we refer readers to Theorem 7.7, Corollary 7.8, and Theorem 7.12.
Theorem 1.4.
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$, and let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Denote by $\xi_{A}$ the characteristic element of $X(A/F_{\infty})$, and let $\mathrm{reg}_{F}$ denote the regular representation of $F$. Assuming Hypothesis $(\mathbf{G})$ holds, we have
In addition, equality holds when Greenberg’s semi-simplicity conjecture 6.2 holds for $A/F$. Moreover, if we assume the finiteness of $\hbox{\cyr X}(A/F)$, then
In fact, the finiteness assumption on the $\ell$-primary part of the Tate-Shafarevich group $\hbox{\cyr X}(A/F)(\ell)$ for a single prime number $\ell$ suffices for the last part of the above theorem by the deep theorem of Kato–Trihan [18].
In order to state our final result, we need to introduce the following stronger finiteness hypothesis. Recall that we write $H={\mathrm{Gal}}(F_{\infty}/F^{\rm cyc})$.
Hypothesis ($\mathbf{H}$). $H^{i}(H,A_{p^{\infty}}(F_{\infty}))$ is finite for all $i\geq 0$.
Let $S$ be a nonempty finite set of primes of $F$ containing all the primes where the abelian variety $A$ has bad reduction, and all the primes of $F$ which ramify in $F_{\infty}$, and let $S^{\prime}$ denotes the subset of primes whose inertia group in $G$ is infinite. We write $A^{*}$ for the dual abelian variety. The following is our final main result.
Theorem 1.5.
Let $A$ be an abelian variety defined over a global field $F$ with $\mathrm{char}F\neq p$ and let $F_{\infty}$ be an admissible $p$-adic Lie extension of $F$. Assume that Hypothesis $(\mathbf{H})$ holds. Then ${\mathrm{Sel}}(A/F_{\infty})$ has finite generalised $G$-Euler characteristic $\chi(G,{\mathrm{Sel}}(A/F_{\infty}))$ if and only if ${\mathrm{Sel}}(A/F^{\rm cyc})$ has finite generalised $\Gamma$-Euler characteristic $\chi(\Gamma,{\mathrm{Sel}}(A/F^{\mathrm{cyc}}))$. Moreover, if this is the case, then
where, for every $v\in S^{\prime}$, $w$ is a fixed prime of $F^{\mathrm{cyc}}$ above $v$.
The above result is a natural analogue of Zerbes [40], where the result was proved in the context of number fields. In the function field context, our result generalises previous computations of Sechi [31] and Valentino [36]. In the case $A$ is an elliptic curve, we may recover the formula of Sechi with Euler factors of the complex $L$-function appearing on the right hand side.
We conclude this introduction with an outline of the paper. In Section 2, we collect some basic notions concerning Iwasawa algebras that will be used throughout the paper. The proof of Theorem 1.3 will be given in Section 3, where the key observation is that the Galois group ${\mathrm{Gal}}(F_{S}/F^{\mathrm{cyc}})$, where $F_{S}$ is the maximal extension of $F$ contained in $\bar{F}$ which is unramified outside $S$, has $p$-cohomological dimension one (see Lemma 3.2). In Section 4, we establish the surjectivity of the localisation map defining the Selmer group ${\mathrm{Sel}}(A/F_{\infty})$ using Jannsen’s spectral sequence and Nekovář’s duality theorem. This will be a key input to our computation of Akashi series in Section 5. In Section 6, we prove a control theorem, showing in particular that the restriction map ${\mathrm{Sel}}(A/F)\to{\mathrm{Sel}}(A/F^{\mathrm{cyc}})^{\Gamma}$ has finite kernel and cokernel. This allows us to relate the order of vanishing at $T=0$ of the characteristic power series of ${\mathrm{Sel}}(A/F^{\mathrm{cyc}})$ to the ${\mathbb{Z}}_{p}$-corank of the Selmer group ${\mathrm{Sel}}(A/F)$. In Section 7, we relate the order of vanishing of characteristic elements evaluated at Artin representations to the Selmer coranks and their twists in the intermediate subextensions of the $p$-adic Lie extension. This, in particular, gives us Theorem 1.4. We then give the definition of the generalised Euler characteristic in Section 8, following the work of Zerbes [40]. Finally, the computations of the generalised Euler characteristics will be carried out in Section 9 under Hypothesis $(\mathbf{H})$ to obtain Theorem 1.5.
2. Prerequisites on Iwasawa algebras
Throughout $p$ will denote a fixed prime. Let $G$ be a compact $p$-adic Lie group without $p$-torsion. The Iwasawa algebra of $G$ over ${\mathbb{Z}}_{p}$ is defined by
[…]
where $U$ runs over the open normal subgroups of $G$ and the inverse limit is taken with respect to the usual projection maps. Suppose for now that $G$ is pro-$p$. Then the ring $\Lambda(G)$ has a skew field of fractions $\mathcal{Q}(G)$ which allows us to define the rank of a $\Lambda(G)$-module $M$ via the following:
The module $M$ is said to be torsion if ${\mathrm{rank}}_{\Lambda(G)}(M)=0$.
Now, suppose that $G$ is not necessarily pro-$p$. By Lazard’s theorem (see [12, Corollary 8.34]), one can always find an open subgroup $G^{\prime}$ of $G$ which is pro-$p$. Then a $\Lambda(G)$-module $M$ is said to be torsion over $\Lambda(G)$ if it is torsion over $\Lambda(G^{\prime})$ in the above sense. We shall also make use of the following equivalent formulation of a torsion $\Lambda(G)$-module: ${\mathrm{Hom}}_{\Lambda(G)}(M,\Lambda(G))=0$.
3. Selmer groups and the ${\mathfrak{M}}_{H}(G)$-Conjecture
Let $p$ be any prime number. Throughout, $F$ will always denote a global function field of characteristic coprime to $p$. We write $k$ for its constant field. Suppose that $A$ is an abelian variety defined over $F$. Let $F_{\infty}$ be a $p$-adic Lie extension of $F$ satisfying the following admissibility conditions:
  1. (i)
  2. (ii)
  3. (iii)
    $F_{\infty}/F$ is unramified outside of a finite set of primes.
@@ 3. Selmer groups and the ${\mathfrak{M}}_{H}(G)$-Conjecture
In particular, $F^{\mathrm{cyc}}$ itself satisfies these conditions. Let $S$ be a nonempty finite set of primes of $F$ containing all the primes where the abelian variety $A$ has bad reduction, and all the primes of $F$ which ramify in $F_{\infty}$.
@@
Let $\bar{F}$ denote a separable closure of $F$. Given any extension ${\mathfrak{K}}$ inside $\bar{F}$, we can define the $p$-primary Selmer group ${\mathrm{Sel}}(A/{\mathfrak{K}})$ of $E$ over ${\mathfrak{K}}$ using the Galois cohomology groups:
@@
where $A_{p^{\infty}}$ denotes the group of all $p$-power division points in $A(\bar{F})$. We note that when ${\mathfrak{K}}$ is an infinite extension of $F$, then ${\mathfrak{K}}_{w}$ is given as the union of completions at $w$ of all finite extensions over $F$ contained in ${\mathfrak{K}}$. If ${\mathfrak{K}}$ is Galois over $F$, the natural action of ${\mathrm{Gal}}({\mathfrak{K}}/F)$ on $H^{1}({\mathfrak{K}},A_{p^{\infty}})$ turns ${\mathrm{Sel}}(A/{\mathfrak{K}})$ into a ${\mathrm{Gal}}({\mathfrak{K}}/F)$-module.
Let $F_{S}$ be the maximal extension of $F$ contained in $\bar{F}$ which is unramified outside $S$. By the choice of $S$, we know $F_{\infty}$ is contained in $F_{S}$. For any extension $K$ of $F$ contained in $F_{S}$, we write $G_{S}(K)$ for the Galois group ${\mathrm{Gal}}(F_{S}/K)$. When $L$ is a finite extension of $F$, we define each prime $v$ of $F$
@@
where $w$ runs over all primes of $L$ lying above $v$. Since $L_{w}$ has characteristic different from $p$, we have $H^{1}(L_{w},A)(p)\simeq H^{1}(L_{w},A_{p^{\infty}})$ by Kummer theory. We shall make this identification without any further mention.
Let $\bar{L}_{w}$ be a separable closure of $L_{w}$, by choosing an embedding $\bar{L}\subset\bar{L}_{w}$, we can view ${\mathrm{Gal}}(\bar{L}_{w}/L_{w})$ as a subgroup of ${\mathrm{Gal}}(\bar{L}/L)$. We then have the following localisation map
(3.2) $\lambda_{S}(A/L):H^{1}(G_{S}(L),A_{p^{\infty}})\to\bigoplus_{v\in S}J_{v}(A/L)$
induced by restriction.
@@
where $L$ runs over all finite extensions of $F$ contained in $K$ and the limit is taken with respect to the restriction maps. We similarly define the localisation map
@@
We will use the following result on Selmer groups in this paper, which can be proven in the same way as in the case of elliptic curves over number fields (see [5, Lemma 2.3]).
@@
Let $K$ be an extension of $F$ contained in $F_{S}$. Then ${\mathrm{Sel}}(A/K)$ satisfies the exact sequence
$0\to{\mathrm{Sel}}(A/K)\to H^{1}(G_{S}(K),A_{p^{\infty}})\stackrel{{\scriptstyle\lambda_{S}(A/K)}}{{\longrightarrow}}\bigoplus_{v\in S}J_{v}(A/K).$
The remainder of the section will be devoted to proving the $\mathfrak{M}_{H}(G)$-conjecture for the Pontryagin dual $X(A/F_{\infty})$ of the $p$-primary Selmer group ${\mathrm{Sel}}(A/F_{\infty})$. In order to achieve this, we start with the following observation which will be crucial in our subsequent discussion.
@@
Since the set $S$ is nonempty and $\mathrm{char}(F)\neq p$, Theorem 10.1.2 (i)(c) from [28] tells us that $\mathrm{cd}_{p}(G_{S}(\bar{k}F))=1$. Note that $G_{S}(\bar{k}F)$ is a closed subgroup of $G_{S}(F^{\mathrm{cyc}})$, and its quotient ${\mathrm{Gal}}(\bar{k}F/F^{\mathrm{cyc}})$ is a profinite group, where every open subgroup has an index prime to $p$. Thus, by applying Proposition 3.3.5 (i) from [28], we find $\mathrm{cd}_{p}(G_{S}(F^{\mathrm{cyc}}))=\mathrm{cd}_{p}(G_{S}(\bar{k}F))$, and the lemma follows. ∎
@@
The cohomology group $H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})$ is cofinitely generated over $\Lambda(H)$. In particular, $H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})^{\vee}$ is a torsion $\Lambda(G)$-module.
@@
Without loss of generality, we may assume that both $G$ and consequently $H$, are pro-$p$ without $p$-torsion. Therefore, by an application of Nakayama’s lemma, we are reduced to showing that $H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})^{H}$ is cofinitely generated over ${\mathbb{Z}}_{p}$. Since $H^{2}(G_{S}(F^{\mathrm{cyc}}),A_{p^{\infty}})=H^{2}(G_{S}(F_{\infty}),A_{p^{\infty}})=0$ by Lemma 3.2, the spectral sequence
@@
degenerates to yield an exact sequence
[…]
@@
Hence, it suffices to show $H^{1}(G_{S}(F_{\mathrm{cyc}}),A_{p})$ is finite, or equivalently, ${\mathrm{corank}}_{{\mathbb{F}}_{p}[[\Gamma]]}\left(H^{1}(G_{S}(F_{\mathrm{cyc}}),A_{p})\right)=0$. Since ${\mathrm{corank}}_{{\mathbb{F}}_{p}[[\Gamma]]}\left(H^{0}(G_{S}(F^{\mathrm{cyc}}),A_{p})\right)=0$ and $H^{2}(G_{S}(F^{{\mathrm{cyc}}}),A_{p})=0$, we have
@@
It then follows from the spectral sequence
$H^{i}\left(\Gamma,H^{j}(G_{S}(F^{\mathrm{cyc}}),A_{p})\right)\Rightarrow H^{i+j}(G_{S}(F),A_{p})$
that the right hand side of (3.4) is
$\sum_{i=0}^{2}(-1)^{i+1}{\mathrm{corank}}_{{\mathbb{F}}_{p}}(H^{i}(G_{S}(F),A_{p})).$
The latter sum vanishes by a global Euler–Poincaré characteristic calculation (cf. [28, (8.7.4)]). This concludes the proof of the lemma. ∎
@@
Let $A$ be an abelian variety defined over a global function field $F$ of characteristic prime to $p$. Let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Then the Selmer group ${\mathrm{Sel}}(A/F_{\infty})$ of $A$ over $F_{\infty}$ is a cofinitely generated $\Lambda(H)$-module. In particular, the $\mathfrak{M}_{H}(G)$-conjecture holds for $X(A/F_{\infty})$.
Furthermore, ${\mathrm{Sel}}(A/F^{\mathrm{cyc}})$ is a cofinitely generated ${\mathbb{Z}}_{p}$-module.
@@
Since ${\mathrm{Sel}}(A/F_{\infty})$ is contained in $H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})$, the conclusion is now an immediate consequence of Lemma 3.3. ∎

### v2

@@ 3. Selmer groups and the ${\mathfrak{M}}_{H}(G)$-Conjecture
In particular, the extension $F^{\mathrm{cyc}}$ itself satisfies these conditions. Let $S$ be a nonempty finite set of primes of $F$ containing all primes where the abelian variety $A$ has bad reduction, as well as all primes of $F$ that ramify in $F_{\infty}$.
@@
Let $\bar{F}$ denote a separable closure of $F$. For any extension $\mathfrak{K}$} inside $\bar{F}$, we define the $p$-primary Selmer group ${\mathrm{Sel}}(A/{\mathfrak{K}})$ using Galois cohomology as follows:
@@
where $A_{p^{\infty}}$ denotes the group of all $p$-power division points in $A(\bar{F})$. Note that when $\mathfrak{K}$} is an infinite extension of $F$, each ${\mathfrak{K}}_{w}$ is defined as the union of completions at $w$ of all finite extensions of $F$ contained in $\mathfrak{K}$}. If $\mathfrak{K}$} is Galois over $F$, then the natural action of ${\mathrm{Gal}}({\mathfrak{K}}/F)$ on $H^{1}({\mathfrak{K}},A_{p^{\infty}})$ endows ${\mathrm{Sel}}(A/{\mathfrak{K}})$ with the structure of a ${\mathrm{Gal}}({\mathfrak{K}}/F)$-module.
Let $F_{S}$ denote the maximal extension of $F$ inside $\bar{F}$ which is unramified outside $S$. By the choice of $S$, we know $F_{\infty}$ is contained in $F_{S}$. For any extension $K$ of $F$ contained in $F_{S}$, we write $G_{S}(K)={\mathrm{Gal}}(F_{S}/K)$. Now, for a finite extension $L/F$ and for each prime $v$ of $F$, define
@@
where the direct sum is taken over all primes $w$ of $L$ lying above $v$. Since $L_{w}$ has characteristic different from $p$, Kummer theory gives the natural identification $H^{1}(L_{w},A)(p)\simeq H^{1}(L_{w},A_{p^{\infty}})$. We shall make use of this identification without any further comment.
Let $\bar{L}_{w}$ be a separable closure of $L_{w}$. By choosing an embedding $\bar{L}\subset\bar{L}_{w}$, we may view ${\mathrm{Gal}}(\bar{L}_{w}/L_{w})$ as a subgroup of ${\mathrm{Gal}}(\bar{L}/L)$. This gives rise to the following localisation map:
(3.2) $\lambda_{S}(A/L):H^{1}(G_{S}(L),A_{p^{\infty}})\to\bigoplus_{v\in S}J_{v}(A/L),$
which is induced by restriction.
@@
where the direct limit is taken over all finite extensions $L$ of $F$ contained in $K$, with respect to the natural restriction maps. Correspondingly, we define the localisation map
@@
We will use the following result on Selmer groups, which can be proved in the same way as for elliptic curves over number fields (see [6, Lemma 2.3]).
@@
Let $K$ be an extension of $F$ contained in $F_{S}$. Then ${\mathrm{Sel}}(A/K)$ fits into the exact sequence:
$0\longrightarrow\mathrm{Sel}(A/K)\longrightarrow H^{1}(G_{S}(K),A_{p^{\infty}})\stackrel{{\scriptstyle\lambda_{S}(A/K)}}{{\longrightarrow}}\bigoplus_{v\in S}J_{v}(A/K).$
The remainder of this section is devoted to proving the $\mathfrak{M}_{H}(G)$-conjecture for the Pontryagin dual $X(A/F_{\infty})$ of the $p$-primary Selmer group ${\mathrm{Sel}}(A/F_{\infty})$. To this end, we begin with a crucial observation.
@@
Since the set $S$ is nonempty and $\mathrm{char}(F)\neq p$, [35, Theorem 10.1.2 (i)(c)] implies that
$\mathrm{cd}_{p}(G_{S}(\bar{k}F))=1.$
Note that $G_{S}(\bar{k}F)$ is a closed subgroup of $G_{S}(F^{\mathrm{cyc}})$, and that its quotient ${\mathrm{Gal}}(\bar{k}F/F^{\mathrm{cyc}})$ is a profinite group in which every open subgroup has an index prime to $p$. Hence, by applying [35, Proposition 3.3.5 (i)], we conclude that $\mathrm{cd}_{p}(G_{S}(F^{\mathrm{cyc}}))=\mathrm{cd}_{p}(G_{S}(\bar{k}F))$, and the lemma follows. ∎
@@
The cohomology group $H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})$ is cofinitely generated over $\Lambda(H)$. In particular, the Pontryagin dual $H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})^{\vee}$ is a torsion $\Lambda(G)$-module.
@@
Without loss of generality, we may assume that both $G$ and consequently $H$, are pro-$p$ and have no $p$-torsion. By Nakayama’s lemma, it suffices to show that $H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})^{H}$ is cofinitely generated over ${\mathbb{Z}}_{p}$.
Since $H^{2}(G_{S}(F^{\mathrm{cyc}}),A_{p^{\infty}})=H^{2}(G_{S}(F_{\infty}),A_{p^{\infty}})=0$ by Lemma 3.2, the Hochschild–Serre spectral sequence
@@
degenerates to yield the following exact sequence
$0\to H^{1}(H,A_{p^{\infty}}(F_{\infty}))\to H^{1}(G_{S}(F^{\mathrm{cyc}}),A_{p^{\infty}})\to H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})^{H}\to H^{2}(H,A_{p^{\infty}}(F_{\infty}))\to 0.$
Since both $H^{1}(H,A_{p^{\infty}}(F_{\infty}))$ and $H^{2}(H,A_{p^{\infty}}(F_{\infty}))$ are cofinitely generated over ${\mathbb{Z}}_{p}$ (see the proof of [18, Theorem 1.1]), it remains to show that $H^{1}(G_{S}(F^{\mathrm{cyc}}),A_{p^{\infty}})$ is cofinitely generated over ${\mathbb{Z}}_{p}$. By Nakayama’s lemma, this is equivalent to showing that $H^{1}(G_{S}(F^{\mathrm{cyc}}),A_{p^{\infty}})[p]$ is finite. From the long exact sequence of Galois cohomology associated with the short exact sequence $0\to A_{p}\to A_{p^{\infty}}\xrightarrow{p}A_{p^{\infty}}\to 0$, we obtain a surjection
@@
Hence, it suffices to show that $H^{1}(G_{S}(F_{\mathrm{cyc}}),A_{p})$ is finite. This is equivalent to proving that
${\mathrm{corank}}_{{\mathbb{F}}_{p}[[\Gamma]]}\left(H^{1}(G_{S}(F_{\mathrm{cyc}}),A_{p})\right)=0.$
Since ${\mathrm{corank}}_{{\mathbb{F}}_{p}[[\Gamma]]}\left(H^{0}(G_{S}(F^{\mathrm{cyc}}),A_{p})\right)=0$ and $H^{2}(G_{S}(F^{{\mathrm{cyc}}}),A_{p})=0$, we have
@@
Using the spectral sequence
$H^{i}\left(\Gamma,H^{j}(G_{S}(F^{\mathrm{cyc}}),A_{p})\right)\Rightarrow H^{i+j}(G_{S}(F),A_{p}),$
we conclude that the right-hand side of (3.3) equals
$\sum_{i=0}^{2}(-1)^{i+1}{\mathrm{corank}}_{{\mathbb{F}}_{p}}(H^{i}(G_{S}(F),A_{p})),$
which vanishes by a global Euler–Poincaré characteristic calculation (cf. [35, (8.7.4)]). This concludes the proof of the lemma. ∎
@@
Let $A$ be an abelian variety defined over a global function field $F$ of characteristic prime to $p$. Let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Then the Selmer group ${\mathrm{Sel}}(A/F_{\infty})$ is a cofinitely generated $\Lambda(H)$-module. In particular, the $\mathfrak{M}_{H}(G)$-conjecture holds for $X(A/F_{\infty})$. Furthermore, ${\mathrm{Sel}}(A/F^{\mathrm{cyc}})$ is a cofinitely generated ${\mathbb{Z}}_{p}$-module.
@@
Since ${\mathrm{Sel}}(A/F_{\infty})$ is contained in $H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})$, the result follows immediately from Lemma 3.3. ∎
We remark that the above theorem relies crucially on the assumption that ${\rm char}(F)\neq p$; otherwise, the key ingredient—the global Euler–Poincaré characteristic—fails, due to the fact that $p=0$ in any ring of characteristic $p$. This also explains the $\mu=0$ condition imposed in [36]. We conclude with the following corollary concerning the generalised $\mu$-invariant, as defined in [18, 46, 4] for finitely generated $\Lambda(G)$-modules.
Corollary 3.5.
The generalised $\mu$-invariant of $X(A/F_{\infty})$ is zero.
Proof.
We have shown that ${\mathrm{Sel}}(A/F_{\infty})$ is a cofinitely generated $\Lambda(H)$-module. By [9, Proposition 2.3], this is equivalent to $X(A/F_{\infty})$ being $\Sigma$-torsion, where $\Sigma$ is the Ore set introduced in Section 7. The result then follows from, for example, the proof of [4, Proposition 3.3 (i)] (also see [18, Lemma 2.7]). ∎